OpenStudy (anonymous):

The expression x - 2 is a factor of p(x) = x^5 - 4x^3 + 2x^2 - 4x + 1. TRUE or FALSE?

jimthompson5910 (jim_thompson5910):

Use the idea that if x - k is a factor of p(x), then p(k) = 0 (this is part of the remainder theorem)

jimthompson5910 (jim_thompson5910):

If you can show that p(2) = 0, then x - 2 is a factor of p(x)

OpenStudy (anonymous):

My answer is False. Am I correct?

jimthompson5910 (jim_thompson5910):

what is p(2)

OpenStudy (anonymous):

I think it's 0?

jimthompson5910 (jim_thompson5910):

p(x) = x^5 - 4x^3 + 2x^2 - 4x + 1 p(2) = (2)^5 - 4(2)^3 + 2(2)^2 - 4(2) + 1 p(2) = ???

OpenStudy (anonymous):

-2

jimthompson5910 (jim_thompson5910):

which calculator are you using?

OpenStudy (anonymous):

a regular calculator.

jimthompson5910 (jim_thompson5910):

graphing calculator? like TI-83?

OpenStudy (anonymous):

no. the regular calculator that comes on a pc

jimthompson5910 (jim_thompson5910):

oh gotcha, I'd use graphcalc or google as a calculator you can type in (2)^5 - 4(2)^3 + 2(2)^2 - 4(2) + 1 exactly as you see it into google

jimthompson5910 (jim_thompson5910):

google will then compute that as a calculator would

jimthompson5910 (jim_thompson5910):

wolfram alpha also does the same

OpenStudy (anonymous):

1

jimthompson5910 (jim_thompson5910):

P(2) = 1 because P(2) does not equal zero, this means x - 2 is not a factor

OpenStudy (anonymous):

I said it was False earlier

jimthompson5910 (jim_thompson5910):

but it's handy to know how/why it's false

jimthompson5910 (jim_thompson5910):

and to know how to compute function values as well

OpenStudy (anonymous):

Thank you .

jimthompson5910 (jim_thompson5910):

you're welcome